OpenStudy (anonymous):

Showing question 6 of 28 Write the series in summation notation: 100 + 95 + 90 + 85 + 80

OpenStudy (michele_laino):

here we have a sequence whose first term is a1=80, and 5-th term is 100, furthermore, the constant of your sequence, is: d=100-95=95-90=90-85=85-80=...

OpenStudy (michele_laino):

furthermore the general term, namely the n-th term is: \[\Large {a_n} = {a_1} + \left( {n - 1} \right)d\]

OpenStudy (michele_laino):

so the requested sum is: \[\Large \sum\limits_{n = 1}^5 {{a_n}} = \sum\limits_{n = 1}^5 {{a_1} + \left( {n - 1} \right)d = ...} \]